# Networked Communication for Decentralised Cooperative Agents in Mean-Field Control

### v1

@@ 4 Theoretical results
To demonstrate the benefits of the networked architecture, we compare it (theoretically here and experimentally in Sec. 5) with the results of modified versions of our algorithm for centralised and independent learners, following the architectures from works learning MFGs online from a non-episodic run of the empirical system (Yardim et al., 2023; Benjamin & Abate, 2023; 2024). In the centralised setting, the Q-network updates of arbitrary agent $i=1$ are automatically pushed to all other agents, and the true global mean-field distribution is always used in place of the local estimate i.e. $\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$. In the independent case, there are no links in $\mathcal{G}^{comm}_{t}$, i.e. $\mathcal{E}^{comm}_{t}=\emptyset$. In this section we prove theoretically that the policy communication and adoption scheme allows networked agents to learn faster than the centralised and independent architectures. Before giving the formal proofs, we suggest informal reasons for our results to aid intuitive understanding.
Remark 2.
@@
Like many cooperative learning paradigms, both the independent and centralised versions of our core learning algorithm (Alg. 2) may suffer from the credit-assignment problem, in that it is not clear how agents’ local state $s^{i}_{t}$ and local action $a^{i}_{t}$ contributed to the (locally estimated) average reward $\tilde{\hat{r}}^{i}_{t}$ (Li & Li, 2024; Cazenille et al., 2025). Agents may receive low individual reward $r^{i}_{t}$ by taking action $a^{i}_{t}$ given $o^{i}_{t}$, but would nevertheless learn that doing so was ‘good’ if the rest of the population took highly rewarded actions at the same step giving high average reward $\tilde{\hat{r}}^{i}_{t}$. By drawing spurious relations, an agent’s updated policy $\pi^{i}_{k+1}(a|o)$ may negatively impact (or simply not advance) the goal of maximising social welfare. While including the (estimated) empirical mean field in the observation $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$ might mitigate this slightly by indicating which mean fields gave high average rewards, this does not solve the issue of allowing learners to distinguish between helpful or unhelpful local actions $a^{i}_{t}$, whether centralised or not. By spreading policies through the population which are expected to have a higher individual return, despite this being a cooperative problem, we reduce the credit-assignment problem by replicating policies that should contribute positively to the population-average return, and filtering out those that do not.
@@
Moreover, even if we assumed credit assignment were not a problem, there is randomness in the Q-network update: agents have stochastic policies and thus may collect a wide variety of transitions to add to their individual buffers, from which they sample randomly when training Q-networks. There may therefore be considerable variance in the quality of their estimated Q-functions, leading in turn to variance in the quality of policy updates. At each iteration of the centralised algorithm, in expectation the central learner will by definition have an average-quality update, and its updated policy will be pushed to the entire population whether or not it performs well. Our decentralised networked approach permits beneficial parallelisation in place of this centralised method, by generating a whole population of possible updates, from which the one(s) estimated to be best-performing can be selected via a process akin to the comparison of fitness functions in evolutionary algorithms.
We prove that networked policy exchange gives faster increase in return than centralised and independent architectures, and give the analysis separately for two important subclasses of cooperative game usually found in MFC, which have different reward structures and therefore require different population behaviour, namely: 1) coordination games, where the social welfare is increased by agents aligning their strategies, such as in consensus/synchronisation/rendezvous tasks; 2) anti-coordination games, where the social welfare is increased by the population exhibiting diverse policies, such as in exploration, coverage or task allocation games. While it is intuitive that adopting policies from neighbours via the communication scheme would be beneficial in coordination games, we find in our experiments that the scheme also benefits populations in anti-coordination games, which we justify theoretically here.
@@
In order to define formally the two types of game, we first introduce the following two functions. $b:\Pi\to\mathbb{R}_{\geq 0}$ is a ‘base return function’ that quantifies a policy’s inherent ability to receive rewards regardless of how many other agents follow the same strategy. [^2] For example, if agents are rewarded for agreeing on one of a number of targets at which to meet, then policies that visit none of the designated targets will have lower returns than those that do, whether agents are aligned or not. $\mathbb{I}[\cdot]$ is the indicator function, which equals 1 if the condition inside is true and 0 otherwise.
Definition 9 (Coordination game).
A coordination game is one where the agents’ return can be decomposed as follows, for all $i,j\in\{1,\dots,N\}$:
$V^{i}(\boldsymbol{\pi},\mu_{0})=h\left(b(\pi^{i}),f_{c}\left(\sum_{j\in\{1,\dots,N\}}\mathbb{I}\left[\pi^{i}=\pi^{j}\right]\right)\right),\;\text{where}$
  • $f_{c}:\mathbb{N}\to\mathbb{R}_{>0}$ is a ‘coordination scaling function’. It has minimum $f_{c}(1)>0$, and increases monotonically with the number of agents whose policies match that of $i$.
@@
  • $h:\mathbb{R}_{\geq 0}\times\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$ is a function that composes $b(\cdot)$ and $f_{c}(\cdot)$ and is monotonic in both arguments, i.e. an increase in either the policy’s intrinsic ability to attain rewards, or the extent to which it is aligned with other agents’ policies, results in a higher return.
Definition 10 (Anti-coordination game).
An anti-coordination game is one where the agents’ return can be decomposed as follows, for all $i,j\in\{1,\dots,N\}$:
$V^{i}(\boldsymbol{\pi},\mu_{0})=h\left(b(\pi^{i}),f_{d}\left(\sum_{j\in\{1,\dots,N\}}\mathbb{I}\left[\pi^{i}=\pi^{j}\right]\right)\right),\;\text{where}$
  • $f_{d}:\mathbb{N}\to\mathbb{R}_{>0}$ is an ‘anti-coordination scaling function’. It has minimum $f_{d}(N)>0$, and increases monotonically with the number of agents whose policies are different from that of $i$.
@@
  • $h:\mathbb{R}_{\geq 0}\times\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$ is a function that composes $b(\cdot)$ and $f_{d}(\cdot)$ and is monotonic in both arguments, i.e. an increase in either the policy’s intrinsic ability to attain rewards, or the extent to which it is different from other agents’ policies, results in a higher return.
For simplicity of the theory, we first make several assumptions that give conditions under which networked agents do outperform centralised ones. The fact that these assumptions do not always hold in practice explains why networked agents may not always outperform centralised ones, though they do in the majority of our experiments.
Recall that at each iteration $k$ of Alg. 2, after independently updating their policies in Line 16, the population has the policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$. There is randomness in these independent policy updates, stemming from the random sampling of each agent’s independently collected buffer. In Lines 1-7 of Alg. 4, agents approximate their infinite discounted individual returns $\{V^{i}(\boldsymbol{\pi},\mu_{0})\}_{i=1}^{N}$ (Def. 2) of their updated policies by computing $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$: the $E$-step discounted return with respect to the empirical mean field generated when agents follow policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$ (i.e. they do not at this stage all follow a single identical policy).
Assumption 1.
@@
Assume that Algs. 1 and 3 allow networked agents to obtain accurate estimations of the true population-average rewards and global empirical mean field respectively, i.e. $\forall i\;\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$ and $\tilde{\hat{r}}^{i}_{t}=\hat{r}$. [^3] In other words, we assume for simplicity that the only difference between the networked and centralised cases is the networked policy communication scheme. In practice, our ablation studies indicate that this is empirically the dominant factor in our experimental settings anyway.
Assumption 2.
Assume that $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ are sufficiently good approximations so as to respect the ordering of the true infinite discounted individual returns $\{V^{i}(\boldsymbol{\pi}_{k+1},\mu_{0})\}_{i=1}^{N}$,i.e. $\forall i,j\in\{1,\dots,N\}$ $\sigma^{i}_{k+1}>\sigma^{j}_{k+1}\;\iff\;V^{i}(\boldsymbol{\pi}_{k+1},\mu_{0})>V^{j}(\boldsymbol{\pi}_{k+1},\mu_{0})$.
Assumption 3.
@@
Assume that directly after the policy updates in Line 16 (Alg. 2), before any policy transfer as in the networked or centralised algorithms, all policies are different (due to the randomness in these updates). This means we have $\min f_{c}$ and $\max f_{d}$. Assume also that each policy has a distinct return, such that $\forall i,j\in\{1,\dots,N\}\;\pi^{i}_{k+1}\neq\pi^{j}_{k+1},V^{i}(\boldsymbol{\pi}_{k+1},\mu_{0})\neq V^{j}(\boldsymbol{\pi}_{k+1},\mu_{0}),\sigma^{i}_{k+1}\neq\sigma^{j}_{k+1}$.
Assumption 4.
Say that $\tau^{comm}_{k}\in\mathbb{R}_{\geq 0}$, such that the softmax adoption scheme (Line 11, Alg. 4) gives non-uniform probabilities of policies being adopted as they are exchanged among neighbours.
Assumption 5.
@@
Assume that after the $C_{p}$ rounds in Lines 8-15 (Alg. 4), in which agents exchange and adopt policies from neighbours, the networked population is left with a single policy such that $\forall i,j\in\{1,\dots,N\}$ $\pi^{i}_{k+1}=\pi^{j}_{k+1}$. [^4] Most simply we can think of Ass. 5 holding if 1) $\tau^{comm}_{k}\rightarrow 0\;\forall k$, such that the softmax essentially becomes a max function, and 2) the communication network $\mathcal{G}^{comm}_{t}$ is static and connected during the $C_{p}$ communication rounds, where $C_{p}$ is equal to the network diameter $d_{\mathcal{G}^{comm}_{t}}$ Under these conditions, previous results on max-consensus algorithms show that all agents in the network will converge on the highest $\sigma^{max}_{k+1}$ value (and hence the unique associated $\pi^{max}_{k+1}$) within a number of rounds equal to the diameter $d_{\mathcal{G}^{comm}_{t}}$ (Nejad et al., 2009; Benjamin & Abate, 2023). However, policy consensus as in Ass. 5 might be achieved even outside of these conditions, including if the network is dynamic and not connected at every step, given appropriate values for $C_{p}$ and $\tau^{comm}_{k}\in\mathbb{R}_{>0}$.
Assumption 6.
@@
We have two different policies that could be shared by the whole population such that $\boldsymbol{\pi}^{x}$ = ($\pi^{x},\dots,\pi^{x}$) and $\boldsymbol{\pi}^{y}$ = ($\pi^{y},\dots,\pi^{y}$). We assume that:
$V^{pop}(\boldsymbol{\pi}^{x},\mu_{0})>V^{pop}(\boldsymbol{\pi}^{y},\mu_{0})\iff W(\pi^{x},I(\pi^{x}))>W(\pi^{y},I(\pi^{y})).$
@@
Call the consensus policy of the networked population $\pi^{\mathrm{net}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{net}}_{k+1}$. Recall that the centralised case is where the Q-network update of arbitrary agent $i=1$ is automatically pushed to all the others instead of the policy evaluation and exchange in Lines 1-15 (Alg. 4); this is equivalent to a networked case where policy consensus is reached on a random one of the policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$. Call this policy arbitrarily given to the whole population $\pi^{\mathrm{cent}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{cent}}_{k+1}$.
Theorem 1.

### v2

@@ 5 Theoretical results
5.1 Introduction
We follow the definitions of the central-agent and independent-learning architectures from closely related works that learn MFGs online from a non-episodic run of the empirical system (policy_mirror_independent; benjamin2024networked; benjamin2024networkedapproximation); both architectures can each be seen as special cases of our networked algorithm:
  • In the central-agent case, only arbitrary central agent $i=1$ updates a Q-network and automatically pushes this to all other agents in place of the decentralised policy communication in Line 17 of Alg. 2. Additionally, the true global mean-field distribution and average reward are always used in place of the local estimates, i.e. $\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$ and $\tilde{\hat{r}}^{i}_{t}=\hat{r}$.
  • In the independent case, there are never any links in $\mathcal{G}^{comm}_{t}$ or $\mathcal{G}^{vis}_{t}$, i.e. $\mathcal{E}^{comm}_{t}=\mathcal{E}^{vis}_{t}=\emptyset$.
We prove theoretically that the policy communication and adoption scheme allows networked agents to increase their returns faster than these alternatives (with the central-agent paradigm being potentially unrealistic and vulnerable in any case). Rem. 5.1 suggests informal reasons for our formal results to aid intuitive understanding.
Remark 5.1.
@@ 5.1 Introduction
Like many cooperative learning paradigms, both the independent and central-agent alternatives to our networked architecture may suffer from the credit-assignment problem, in that it is not clear how agents’ local state $s^{i}_{t}$ and local action $a^{i}_{t}$ contributed to the (locally estimated) average reward $\tilde{\hat{r}}^{i}_{t}$ (10777009; cazenille2025signalling). Agents may receive low individual reward $r^{i}_{t}$ by taking action $a^{i}_{t}$ given $o^{i}_{t}$, but would nevertheless learn that doing so was ‘good’ if the rest of the population took highly rewarded actions at the same step giving high average reward $\tilde{\hat{r}}^{i}_{t}$. By drawing spurious relations, an agent’s updated policy $\pi^{i}_{k+1}(a|o)$ may negatively impact (or simply not advance) the goal of maximising social welfare. Including the (estimated) empirical mean field in the observation $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$ might mitigate this slightly by indicating which mean fields gave high average rewards. However, this does not solve the issue of allowing learners to distinguish between helpful or unhelpful local actions $a^{i}_{t}$, whether those learners are centralised or not, since actions can affect rewards in ways other than simply by helping to reach a certain mean field. By updating policies with respect to average return but then spreading updates through the population which are estimated to give a higher individual return, despite this being a cooperative problem, we reduce the credit-assignment problem by replicating updated policies that should contribute positively to the population-average return, and filtering out those that do not.
@@
Moreover, even if we assumed credit assignment were not a problem, there is randomness in the Q-network update: agents have stochastic policies and thus may collect a wide variety of transitions to add to their individual buffers, from which they sample randomly when training Q-networks. There may therefore be considerable variance in the quality of their estimated Q-functions, leading in turn to variance in the quality of policy updates. At each iteration of the central-agent algorithm, in expectation the central learner will by definition have an average-quality update, and its updated policy will be pushed to the entire population whether or not it performs well, also giving large variance in the quality of updates. Our decentralised networked approach permits beneficial parallelisation in place of this single-learner method, by generating a whole population of possible updates, from which the one(s) estimated to be best-performing can be selected via a process akin to the comparison of fitness functions in evolutionary algorithms. These are then spread around the population, biasing networked populations towards better performing updates.
We give the theoretical analysis separately for two important subclasses of cooperative game usually found in MFC, which have different reward structures and therefore can incentivise different population behaviour:
  1. 1.
    coordination games, where the social welfare is increased by agents aligning their strategies, such as in consensus/synchronisation/rendezvous tasks;
  2. 2.
    anti-coordination games, where the social welfare is increased by the population exhibiting diverse strategies, such as in exploration, coverage or task allocation games.
These subclasses cover a large proportion of cooperative objectives in anonymous, symmetric settings with large populations. We emphasise that the fact that agents would in principle benefit from having diverse policies in anti-coordination games does not contradict the classical MFC framework that simplifies the infinite population problem by finding the single policy to be shared by all agents. In the symmetric (i.e. identical reward and transition functions) MFC limit, an optimal solution can be realised by having the infinite agents all follow the single socially optimal policy, even for reward functions that favour diversity. A very large number of works on both MFC and MFGs conduct experiments on anti-coordination games, particularly dispersal and exploration tasks, despite assuming that the population follows a shared single policy learnt by a central node (doi:10.1073/pnas.1922204117; survey_learningMFGs; lee2024mean). We make the distinction between coordination and anti-coordination games to aid theoretical analysis of our decentralised policy adoption scheme compared with entirely independent learning: while it is intuitive that adopting independently-updated policies from neighbours via the communication scheme would be beneficial in coordination games, we also show theoretically and empirically that the adoption scheme provides a benefit in anti-coordination games, though this requires separate analysis.
@@
To define the two types of game, we first introduce the following functions. $\mathbb{I}[\cdot]$ is the indicator function, which equals 1 if the condition inside is true and 0 otherwise. $b:\Pi\to\mathbb{R}_{\geq 0}$ is a base return function that quantifies a policy’s inherent ability to receive rewards regardless of how many other agents follow the same strategy. For example, if agents are rewarded for agreeing on one of a number of targets at which to meet, then policies that visit none of the designated targets will have lower returns than those that do, whether agents are aligned or not. $f_{c}:\mathbb{N}\to\mathbb{R}_{>0}$ (resp. $f_{d}:\mathbb{N}\to\mathbb{R}_{>0}$) is a coordination (resp. anti-coordination) scaling function. It has minimum $f_{c}(1)>0$ (resp. $f_{d}(0)>0$), and increases monotonically with the number of agents whose policies match (resp. are different from) $i$’s.
Definition 5.2 (Coordination game).
@@
The agents’ return can be decomposed as follows, $\forall i,j\in\{1,\dots,N\}$: $V^{i}(\bm{\pi},\mu_{\bar{t}})=h\left(b(\pi^{i}),f_{c}\left(\sum_{j\in\{1,\dots,N\}}\mathbb{I}\left[\pi^{i}=\pi^{j}\right]\right)\right)$, where $h:\mathbb{R}_{\geq 0}\times\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$ is a function that composes $b(\cdot)$ and $f_{c}(\cdot)$ and is monotonic in both arguments, i.e. an increase in either the policy’s intrinsic ability to attain rewards, or the extent to which it is aligned with other agents’ policies, gives a higher return.
Definition 5.3 (Anti-coordination game).
@@
The agents’ return can be decomposed as follows, $\forall i,j\in\{1,\dots,N\}$: $V^{i}(\bm{\pi},\mu_{\bar{t}})=h\left(b(\pi^{i}),f_{d}\left(N-\sum_{j\in\{1,\dots,N\}}\mathbb{I}\left[\pi^{i}=\pi^{j}\right]\right)\right)$, where $h:\mathbb{R}_{\geq 0}\times\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$ is a function that composes $b(\cdot)$ and $f_{d}(\cdot)$ and is monotonic in both arguments, i.e. an increase in either the policy’s intrinsic ability to attain rewards, or the extent to which it is different from other agents’ policies, gives a higher return.
Note that in our setting, where policy parameters are directly communicated and adopted among the population, we focus on exact equality of policies for simplicity of the theory. However, these definitions could be made more general and inclusive by instead considering similarity kernels or label mappings of strategically relevant parts of policies.
5.2 Analysis
Our sub-routines involve time-varying networks sharing different types of information at different points in the algorithm, meaning that theoretical analysis can potentially grow complicated. We seek to simplify this analysis to make it more intuitive and useful by focusing on the benefit of the decentralised policy exchange scheme in Alg. 3. This is because our ablation studies of Algs. 1 (average reward estimation) and 4 (mean field estimation) in Sec. 6.3 indicate that the policy exchange scheme is the dominant factor in driving the benefit of the networked paradigm in our experimental settings. Moreover, recall that Alg. 4 is only necessary when we allow population-dependent policies i.e. $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$, whereas for stationary problems, including all those in our experiments and many others, using a mean field observation or estimation is not actually required for finding the optimal policy.
For simplicity of the theory, we make several assumptions. We explore the conditions under which these assumptions apply in practice, and discuss how even when loosening the assumptions, they still provide useful heuristic insight as to how our networked communication scheme affords benefits over the central-agent and independent-learning architectures. We do not enforce the assumptions in our experiments, and our empirical results nevertheless follow our theoretical theorems in all but some specific instances that we discuss.
The first assumption simplifies the theory by presuming that it is only the decentralised policy communication scheme that creates a difference in learning between the networked and central-agent cases, by assuming that the estimated mean fields and average rewards are equivalent to the true ones used in the central-agent case. Note that populations with fully connected networks will in any case always be able to accurately estimate $\hat{r}$ and ${\hat{\mu}}_{t}$ by Algs. 1 and 4, even for $C_{r}=1$ and $C_{e}=0$. This is may apply reasonably commonly in practice depending on the scenario; for example, if the network is defined by a broadcast radius (as in our experiments), then the network will be fully connected whenever that radius is at least large enough to cover the area that all the agents fall within. Moreover, as just mentioned, our ablation studies suggest that the policy communication scheme is the dominant factor in our experimental settings anyway, with the estimated mean field not required at all in the broad class of stationary problems. We leave analysis of the theoretical impact of worsening mean-field and average-reward estimations for future work.
Assumption 5.4.
@@ 5.2 Analysis
Assume that Algs. 1 and 4 allow networked agents to obtain accurate estimations of the true population-average rewards and global empirical mean field respectively, i.e. $\forall i\;\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$ and $\tilde{\hat{r}}^{i}_{t}=\hat{r}$.
Recall that at each iteration $k$ of Alg. 2, after individually updating their policies in Line 16, the population has the policies $\pi^{i}_{k+1}$$_{i=1}^{N}$. There is randomness in these individual policy updates, stemming from the random sampling of each agent’s individually collected buffer. In Lines 1-7 of Alg. 3, agents estimate the individual infinite-step discounted returns $\{V^{i}(\bm{\pi},\mu_{0})\}_{i=1}^{N}$ (Def. 3.2) of their updated policies by computing $\sigma^{i}_{k+1}$$_{i=1}^{N}$: the $E$-step discounted return with respect to the empirical mean field generated when agents follow policies $\pi^{i}_{k+1}$$_{i=1}^{N}$.
We next assume that the populations’ policies are all pair-wise distinct after the updates in Line 16 and before the policy communication. This ensures that policies that are estimated to receive higher returns (and are thus adopted) are being evaluated as higher-performing due to receiving higher base returns, rather than simply because of how aligned or distinct they already happen to be with regard to other policies. This avoids scenarios where, for example, significantly suboptimal policies that are shared across multiple agents after the update (in the case of a coordination game) end up spreading through the population by communication at the expense of a more promising but less common policy, decelerating rather than accelerating improvement. In practice, this assumption is highly likely to apply in most situations in any case. Even if agents start a given iteration with identical policies, their different random seeds are likely to mean that they collect different sample transitions to add to their reinitialised buffers. Even if their buffers end up containing identical transitions, their different random seeds are likely to mean that they sample differently from their buffers, leading to slightly different updates to their policy networks.
Assumption 5.5.
@@
Assume that directly after the policy updates in Line 16 (Alg. 2), before any policy transfer as in the networked or central-agent algorithms, all policies are pair-wise distinct due to the randomness in these updates, i.e. $\forall i,j\in\{1,\dots,N\}\;\pi^{i}_{k+1}\neq\pi^{j}_{k+1}$. This means the function $f_{c}$ attains its minimum $f_{c}(1)$, and $f_{d}$ attains its maximum $f_{d}(N-1)$.
We now assume that the finite-step estimations of the returns give sufficiently accurate comparisons between policies, so that better policies are indeed the ones that get adopted in expectation.
Assumption 5.6.
Assume that $\sigma^{i}_{k+1}$$_{i=1}^{N}$ are sufficiently good estimations so as to respect the ordering of the true infinite discounted individual returns $\{V^{i}(\bm{\pi}_{k+1},\mu_{0})\}_{i=1}^{N}$, i.e.
$V^{i}(\bm{\pi}_{k+1},\mu_{0})>V^{j}(\bm{\pi}_{k+1},\mu_{0})\;\;\iff\;\;\sigma^{i}_{k+1}>\sigma^{j}_{k+1}\;\;\;\;\;\;\;\;\;\;\forall i,j\in\{1,\dots,N\}.$
In practice, even if Assumption 5.6 does not strictly hold, the softmax parameter $\tau^{comm}_{k}$ allows a smooth degradation as the ordering of the approximations worsens with respect to the ordering of the true values. That is, if instead of the exact correct policy ordering we have that better policies are simply more likely to be given higher estimated evaluations, then the softmax means that these policies remain more likely to spread, and a better policy may still be adopted even if it is not evaluated as being better.
The next assumption presumes that the networked population reaches consensus on a single policy within each $k$ iteration. We use it only in Thm. 5.9, and we do so to give general and intuitive comparison with the central-agent population which always shares a single policy. Incomplete consensus would give different levels of alignment/diversity, such that the relative performance of the central-agent and networked architectures might otherwise depend on the specific reward function of the task, and whether base return or alignment/diversity is more important in that reward function.
Assumption 5.7.
Assume that after the $C_{p}$ rounds in Lines 8-15 (Alg. 3), in which agents exchange and adopt policies from neighbours, the networked population is left with a single policy such that $\forall i,j\in\{1,\dots,N\}$ $\pi^{i}_{k+1}=\pi^{j}_{k+1}$.
While this may sound like a strong assumption, we phrase it like this so as not to make overly strong restrictions on the communication network instead - we intentionally leave it so that Assumption 5.7 can be fulfilled in numerous ways. Most simply we can think of Assumption 5.7 holding if:
  1. 1.
    we set $\tau^{comm}_{k}$ close to 0 for all $k$, such that the softmax essentially becomes a max function; and
  2. 2.
    the communication network $\mathcal{G}^{comm}_{t}$ is static and connected during the $C_{p}$ communication rounds, where $C_{p}$ is at least as large as the network diameter $d_{\mathcal{G}^{comm}_{t}}$.
@@
Under these conditions, previous results on max-consensus algorithms show that all agents in the network will converge on the highest value $\sigma^{max}_{k+1}$ (and hence the unique associated $\pi^{max}_{k+1}$) within a number of rounds equal to the diameter $d_{\mathcal{G}^{comm}_{t}}$ (5348437; benjamin2024networked). If we assumed more strongly that the network was always fully connected, policy consensus would be achieved within a single communication round.
Policy consensus can be achieved even outside of these conditions, including if the network is dynamic and not connected at every step. The union of a collection of graphs $\mathcal{G}_{t},\mathcal{G}_{t+1},\cdots,\mathcal{G}_{t+\omega}\}$ ($\omega\in\mathbb{N}$) is the graph with vertices and edge set equalling the union of the vertices and edge sets of the graphs in the collection (1205192). A collection is jointly connected if its members’ union is connected. Now, instead of assuming that the communication network is static and connected, we assume instead only that the sequence of networks contains one or more jointly connected collections. Then max-consensus is reached within $C_{p}$ if $C_{p}$ is large enough that the number of jointly connected collections occurring within $C_{p}$ is equal to the largest diameter of the union of any such collection.
Thus Assumption 5.7 may not hold if $C_{p}$ is not large enough or if parts of the population remain isolated. However, we do not enforce this assumption in our experiments, where we use $C_{p}=1$ to show the benefit of even just one communication round, yet we still see networked populations significantly outperforming central-agent populations across anti-coordination games. In coordination games, while networked populations that are more connected (due to having larger communication radii) usually perform similarly to or better than central-agent populations, those that are less connected occasionally perform less well than the central-agent populations. This is probably due to Assumption 5.7 being empirically more likely to be violated in less connected populations, which in turn is more of an issue in coordination games (where consensus is more likely to be beneficial) than in anti-coordination games (where some lack of consensus does not prevent, or even helps, networked populations to outperform central-agent ones in practice).
The next assumption presumes that if a certain policy, when followed by all members of a finite population, is better than another policy when the latter is followed by all members of a finite population, then the same quality ordering will apply when members of infinite populations follow each policy. We require this in order to relate our analysis of learning in the empirical finite population back to the mean field limit when comparing with central-agent learning. Since the finite population can be arbitrarily large, and in many environments when all agents follow the same policy the finite population-average return will converge smoothly to the infinite population social welfare, this assumption will naturally hold in many scenarios. For example, a policy that is better than another at getting a population of 500 or 5,000,000 agents to cluster in a particular location will also be better than the other policy at getting an infinite population to gather at the location. Nevertheless this order preservation is not a completely general phenomenon, and strict inequalities can vanish or reverse in the limit, especially in models with thresholds or discontinuities in the dependence of rewards or transitions on the mean field, so we state it as an explicit condition.
Assumption 5.8.
@@
Say we have two different policies that could be shared by the whole population such that $\bm{\pi}^{x}$ = ($\pi^{x},\dots,\pi^{x}$) and $\bm{\pi}^{y}$ = ($\pi^{y},\dots,\pi^{y}$). We assume that:
$V^{pop}(\bm{\pi}^{x},\mu_{0})>V^{pop}(\bm{\pi}^{y},\mu_{0})\iff W(\pi^{x},I(\pi^{x}))>W(\pi^{y},I(\pi^{y})).$
@@
We have now given all the assumptions for our first theorem. Assumption 5.7 assumes that after the $C_{p}$ policy exchange rounds in Lines 8-15 of Alg. 3, the networked population is left with a single policy. Call this consensus policy $\pi^{\mathrm{net}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{net}}_{k+1}$. Recall that the central-agent case is where the Q-network update of arbitrary agent $i=1$ is automatically pushed to all the others instead of the policy evaluation and exchange in Line 17 of Alg. 2; this is equivalent to a networked case where policy consensus is reached on a random one of the policies $\pi^{i}_{k+1}$$_{i=1}^{N}$. Call this policy arbitrarily given to the whole population $\pi^{\mathrm{cent}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{cent}}_{k+1}$.
We can now give our first theorem, namely that in expectation networked populations will increase their returns at least as fast as central-agent ones.
Theorem 5.9.